Consider two neighboring island countries called Bellissima and Felicidad. They each have 4 million labor hours available perIn the following table, enter each countrys production decision on the third row of the table (labeled Production). Bellis
Answer #1

Blanks-

1) 16/8 = 2

2) 20/5 = 4

3) Bellissima

4) Felicidad

5) 8*4 = 32

6) 20*4 = 80

7) 32-23 = 9

8) 80-68 = 12

Bellissima

Felicidad

Corn

Jeans

Corn

Jeans

Without Trade

Production

8

48 15 20

Consumption

8

48 15 20

With Trade

Production

32

0 0 80

Trade action

Export 18

Import 54

Import 18

Export 54

Consumption

14

54

18

26

Gains from Trade

Increase in Consumption

6

6 3 6